# Spatial Analysis for AI-segmented Histopathology Images: Methods and Implementation

### v1

@@ 4.1 Implementation
SASHIMI is a browser-based computational framework implemented in R and Python for the extraction, visualization, and computation of spatial features from AI-segmented histopathology images, enabling real-time analysis. The platform functions as both an exploratory and feature-extraction tool, generating two complementary forms of output from marked point pattern data: (i) functional visualizations, which characterize distance-dependent spatial dynamics (e.g., Ripley’s $K$-function and pair correlation function), and (ii) scalar indices, which quantify spatial autocorrelation and compositional similarity across tissue regions (e.g., Moran’s I, Jaccard index, and summary statistics derived from persistence diagrams). Together, these outputs provide a comprehensive and interpretable representation of the spatial organization and topological structure of tumor tissue architecture.
@@ 4.2 Input Requirements and Data Format
SASHIMI accepts tabular data in CSV format with a maximum file size of $4$ MB. Each input file must contain three columns corresponding to the $x$-coordinate, $y$-coordinate, and cell type. Upon upload, the data are displayed as a point cloud, where each point represents an individual cell on the tissue slide and is color-coded by cell type. Users may select up to three cell types for analysis. For each selection, SASHIMI computes both single-type spatial features and pairwise cross-type spatial functions.
@@ 4.3 Output Types and Visualization
SASHIMI produces two forms of output. Functional outputs consist of spatial summary statistics rendered as plots, enabling interpretation of distance-based spatial relationships. Scalar outputs include autocorrelation measures, similarity indices, and summary statistics of the persistence diagram. Returned in tabular form, it provides quantitative measures suitable for direct comparison and statistical analysis. All computed features can be downloaded in CSV format for downstream use.
@@ 5 Case Studies
To examine the relationship between cellular spatial organization and patient survival outcomes, we analyzed two real-world datasets: lung cancer pathology images from NSCLC patients in the NLST and oral tissue pathology images from OPMD patients in the EPOC trial at The University of Texas MD Anderson Cancer Center.
To evaluate the prognostic relevance of spatial features, functional Cox proportional hazards (FCoxPH) models were fitted using the survival package in R. Patient ID was used as a clustering variable to account for within-patient correlations arising from multiple image regions. Each spatial feature was modeled individually, with patient age and cancer stage included as additional scalar covariates. Statistical significance was assessed across multiple thresholds to identify spatial descriptors potentially associated with patient prognosis.
To characterize distance-dependent spatial dynamics, we applied functional principal component analysis (FPCA) to the computed functional descriptors. Spatial functions were decomposed using FPCA based on the Karhunen–Loève expansion (kong2018flcrm), and the first two principal components, which captured approximately 90% of the total variance, were retained. The resulting FPCA scores were then used as covariates in the survival analysis models, along with additional scalar covariates. Additional methodological details are provided in the supplementary materials.
@@ 5.1 Case Study I: Lung Cancer Pathology Images
Collectively, these findings demonstrate that both conventional spatial statistics and topological descriptors provide complementary insights into the spatial organization of cells. In particular, topological features capture higher-order structural information not evident from distance-based measures alone.
@@ 5.2 Case Study II: Oral Tissue Pathology Images
The OPMD dataset consisted of $255$ whole-slide images acquired at $40\times$ magnification from 128 patients with oral premalignant lesions. Using a deep convolutional neural network developed by our team (wang2019convpath), nuclei were segmented and classified into four distinct categories: basal, other epithelial, lymphocytes, and stromal cells. For feature computation, only three types—basal, lymphocyte, and stromal cells—were used, excluding the other epithelial category. This resulted in $703$ segmented image regions, each containing approximately $1000$ cell points.
@@
Overall, the OPMD dataset exhibited a distinct pattern of significant spatial predictors compared with the NLST cohort, with $J$-functions and topological feature counts playing a prominent role in survival prediction.
@@
5.3 Threshold-based Significance Analysis
In the NLST dataset, $16$ spatial features reached significance at $p<0.20$, with $11$ features remaining significant at $p<0.15$, 77 features at $p<0.10$, and $3$ features at $p<0.05$. Similarly, the OPMD dataset demonstrated $15$ significant features at $p<0.20$, $10$ features at $p<0.15$, 88 features at $p<0.10$, and $3$ features achieving significance at $p<0.05$. This gradient of statistical significance suggests that spatial organization encodes meaningful prognostic information, although the effect sizes may be subtle due to the inherent variability in cellular spatial arrangements within tumor tissue architecture and other domain knowledges such as tumor size in different types of cancers.
@@ 5.3 Threshold-based Significance Analysis
The use of multiple significance thresholds is justified by several factors inherent to spatial point pattern analysis. First, the resolution and density of marked point pattern data can influence the sensitivity of spatial statistics to detect biologically relevant structures. Second, the complex multi-scale nature of tumor spatial organization may yield moderate rather than strong statistical associations. Third, the exploratory nature of comprehensive spatial feature analysis warrants consideration of features that may achieve stronger significance in larger cohorts or independent validation datasets.
@@ 8 Data Availability Statement
Case study II. The OPMD dataset analyzed in this study is available in the Zenodo repository under restricted access at https://doi.org/10.5281/ZENODO.10664056. This dataset is described in  (piyarathne2024comprehensive). The nuclei segmentation and classification results generated in this study using ConvPath (wang2019convpath) are available from the corresponding author upon reasonable request.

### v2

@@ 4.1 Implementation
SASHIMI is a browser-based computational framework implemented in R and Python for the extraction, visualization, and computation of spatial features from AI-segmented histopathology images, enabling real-time analysis. The web-based design is intentionally practitioner-facing: analyses are executed through the browser interface without requiring end-users to install a programming compiler, manage external package dependencies, or configure local environments. This deployment choice directly addresses common barriers for translational users who wish to interrogate AI-segmented pathology outputs but lack specialized computational infrastructure. Consistent with prior work on practitioner-oriented post-segmentation analytics (Fernández et al., 2022), SASHIMI treats AI-segmented results as the analysis-ready input and focuses on producing interpretable spatial and topological descriptors that can be used immediately for exploratory investigation or downstream modeling. The platform functions as both an exploratory and feature-extraction tool, generating two complementary forms of output from marked point pattern data: (i) functional visualizations, which characterize distance-dependent spatial dynamics (e.g., Ripley’s $K$-function and pair correlation function), and (ii) scalar indices, which quantify spatial autocorrelation and compositional similarity across tissue regions (e.g., Moran’s I, Jaccard index, and summary statistics derived from persistence diagrams). Together, these outputs provide a comprehensive and interpretable representation of the spatial organization.
@@ 4.2 Input Requirements and Data Format
SASHIMI accepts tabular data in CSV format with a maximum file size of $4$ MB. Inputs are assumed to be derived from upstream AI-based segmentation pipelines (i.e., preprocessed cell detections and type annotations), so SASHIMI does not operate on raw whole-slide images; rather, it consumes segmentation-derived cell-level coordinates and labels as a standardized interchange format for post-segmentation spatial analysis. Each input file must contain three columns corresponding to the $x$-coordinate, $y$-coordinate, and cell type. Upon upload, the data are displayed as a point cloud, where each point represents an individual cell on the tissue slide and is color-coded by cell type. Users may select up to three cell types for analysis. For each selection, SASHIMI computes both single-type spatial features and pairwise Cross-type spatial functions.
@@ 4.3 Output Types and Visualization
SASHIMI produces two forms of output. Functional outputs consist of spatial summary statistics rendered as plots, enabling interpretation of distance-based spatial relationships. Scalar outputs include autocorrelation measures, similarity indices, and summary statistics of the persistence diagram. Returned in tabular form, it provides quantitative measures suitable for direct comparison and statistical analysis. All computed features can be downloaded in CSV format for downstream use. By coupling immediate visualization with downloadable, analysis-ready feature tables, SASHIMI is designed to support rapid exploratory workflows for practitioners—allowing users to screen spatial patterns, generate hypotheses, and export standardized outputs for subsequent statistical modeling without additional software setup on the user side.
@@ 5 Case Studies
To illustrate the use of SASHIMI as an exploratory spatial analysis and feature extraction platform, we analyzed two real-world datasets: lung cancer pathology images from NSCLC patients in the NLST and oral tissue pathology images from OPMD patients in the EPOC trial at The University of Texas MD Anderson Cancer Center.
Rather than performing a comprehensive evaluation or comparative assessment of spatial features, these case studies are intended to demonstrate how SASHIMI facilitates systematic extraction, visualization, and organization of diverse spatial descriptors from AI-segmented histopathology images. To provide a concrete illustration of downstream analysis, functional Cox proportional hazards (FCoxPH) models were fitted using the survival package in R. Patient ID was used as a clustering variable to account for within-patient correlations arising from multiple image regions. Each spatial feature was modeled individually, with patient age and cancer stage included as additional scalar covariates. Statistical significance was assessed across multiple thresholds in an exploratory manner, with the goal of highlighting representative spatial patterns that may warrant further investigation rather than establishing definitive prognostic biomarkers. To characterize distance-dependent spatial dynamics, we applied functional principal component analysis (FPCA) to the computed functional descriptors. Spatial functions were decomposed using FPCA based on the Karhunen–Loève expansion (Kong et al., 2018), and the first two principal components, which captured approximately 90% of the total variance, were retained. The resulting FPCA scores were then used as covariates in the survival analysis models, along with additional scalar covariates. Additional methodological details are provided in the supplementary materials.
@@ 5.1 Case Study I: Lung Cancer Pathology Images
Collectively, these findings demonstrate that conventional spatial statistics and topological descriptors provide complementary insights into the spatial organization of the tumor microenvironment. Notably, Cross-type interactions consistently emerged as the most prognostic features, outperforming single-type or type-agnostic measures. This pattern underscores the importance of explicitly incorporating cell-type information when characterizing spatial proximity to tumor cells, aligning with established biological principles regarding the functional role of intercellular interactions in cancer progression  (Hanahan and Weinberg, 2011). In particular, immune–tumor and immune–stromal interactions exhibited the strongest statistical associations with survival outcomes, whereas stromal–tumor relationships showed weak or negligible correlations with patient prognosis.
When examined by feature class, topological descriptors derived from the witness complex and pair correlation functions achieved statistical significance at the most stringent threshold. Although both the pair correlation function and the K-function quantifies distance-dependent clustering, the K-function failed to capture meaningful prognostic signals in this setting. This limitation arises from its cumulative formulation, which aggregates the number of neighboring cells within an increasing radius r, thereby emphasizing overall abundance rather than local spatial dynamics. In contrast, the pair correlation function evaluates cell interactions at fixed spatial scales, isolating distance-specific deviations from spatial randomness. This distinction enabled the detection of prognostically relevant immune–tumor spatial relationships, reinforcing the importance of spatial proximity over coarse measures of cellular abundance within the tumor microenvironment (Shen et al., 2024; Backman et al., 2023; Yuan, 2016).
@@ 5.2 Case Study II: Oral Tissue Pathology Images
The OPMD dataset consisted of $255$ whole-slide images acquired at $40\times$ magnification from 128 patients with oral premalignant lesions. Using a deep convolutional neural network developed by our team (Wang et al., 2019), nuclei were segmented and classified into four distinct categories: basal, other epithelial, lymphocytes, and stromal cells. For feature computation, only three types—basal, lymphocyte, and stromal cells—were used, excluding the other epithelial category. This resulted in $703$ segmented image regions, each containing approximately $1000$ cell points.
@@
Overall, the OPMD dataset exhibited a distinct pattern of significant spatial predictors compared with the NLST cohort, with J-functions and topological feature counts playing a particularly prominent role in survival prediction. In contrast to the NLST dataset, both Cross-type and single-type interactions contributed substantially to prognostic modeling in OPMD. Among these, tumor–stromal and tumor–immune interactions demonstrated the strongest associations with survival outcomes, followed by tumor-only interaction patterns. Notably, several single-type and aggregation-based spatial functions emerged as significant predictors in OPMD, despite showing little or no prognostic relevance in the NLST cohort. In contrast, topological features consistently retained their predictive value across both datasets, underscoring their robustness to cohort-specific variation.
Given the smaller segmented batch size of the OPMD dataset—reflecting inherent differences in lesion size, with approximately 1,000 cell points per image compared to an average of 12,000 in the NLST dataset—the enhanced performance of aggregation-based functions such as the J-function is likely attributable to their sensitivity to local spatial structure. This suggests that, under constrained image sizes, spatial summary functions may preferentially capture fine-scale organization that becomes diluted in larger fields of view. Consequently, these findings highlight the batch-size sensitivity of certain spatial descriptors and emphasize the importance of considering image scale and sampling density when interpreting spatial feature performance across heterogeneous cohorts.
@@
5.3 Discussion
A key advantage of integrating multiple classes of spatial descriptors within SASHIMI is the ability to characterize complementary aspects of cellular organization across spatial scales. In this section, we aim to provide a detailed overview of the suite of features by discussing their respective capabilities and limitations.
Radial distribution functions, such as the pair correlation function and mark connection function, emphasize local spatial interactions by quantifying deviations from spatial randomness at specific distance ranges. These features are particularly sensitive to pairwise cellular proximity and localized patterns of structure or avoidance, which are frequently considered biologically relevant in tumor–immune interactions. In contrast, other members of the spatial summary statistics family—namely cumulative descriptors such as Ripley’s K-, G-, and L-functions—aggregate spatial information over increasing radii, capturing more global or multi-scale trends in cellular aggregation. While these cumulative summaries provide a broad overview of spatial organization, they may dilute fine-scale interaction signals when cell density is high or when spatial heterogeneity varies across regions or cohorts. When applied to the two real datasets, radial distribution functions demonstrated stronger correlations with survival outcomes in the NLST dataset compared to the OPMD dataset, where the average number of cells per image batch is substantially larger. Conversely, aggregation functions showed stronger correlations in the OPMD dataset and more moderate associations in NLST, suggesting that prognostic signals in OPMD may be driven by local proximity structure of cells rather than overall abundance or composition (Shen et al., 2024; Backman et al., 2023; Yuan, 2016).
While the inclusion of both distance-specific and cumulative descriptors within SASHIMI allows users to examine spatial organization from complementary local and global perspectives without presupposing a single “optimal” summary statistic, the scope of spatial structure captured by each feature remains dependent on the choice of image batch size.
Topological descriptors derived from persistent homology offer a distinct perspective on spatial organization by capturing connectivity and structural patterns that are invariant to continuous deformations and less sensitive to local fluctuations in cell density. Unlike distance-based summaries, which depend explicitly on spatial scale parameters and image size, topological features summarize the birth, persistence, and disappearance of connected components and loops across multiple spatial resolutions. In the context of tumor tissue architecture, these properties make topological features well suited for capturing higher-order structural organization, such as compartmentalization and multiscale connectivity between tumor, stromal, and immune regions. Their consistent emergence as some of the strongest covariates across datasets in the illustrative case studies suggests that topological summaries may provide robust, complementary information beyond classical proximity-based measures, particularly in heterogeneous or sparsely sampled imaging settings.
Lastly, the areal feature family exhibited a moderate level of correlation in the OPMD dataset but weak or no meaningful association in the NLST dataset. Similar to aggregation-based functions, areal features computed on a fixed grid resolution limit the precise capture of fine local structure within the tumor microenvironment, as they primarily reflect compositional similarity and regional overlap based on cell abundance.
Another recurring pattern observed across the case studies is the frequent prominence of Cross-type spatial descriptors relative to single-type summaries. Cross-type features explicitly quantify spatial relationships between distinct cell populations, such as tumor–immune or tumor–stromal interactions, thereby encoding biologically meaningful aspects of the tumor microenvironment that cannot be inferred from single-population distributions alone, further recognizing the strong prognostic relevance of cell-type-specific spatial organization.
By systematically organizing both single-type and Cross-type spatial descriptors, SASHIMI enables users to explore these interaction-driven patterns in a unified and interpretable manner, facilitating hypothesis generation regarding the spatial mechanisms underlying disease progression. Importantly, SASHIMI is not designed to rank or prioritize spatial features based on predictive performance. Instead, its primary goal is to provide a standardized and accessible interface for computing, visualizing, and organizing diverse spatial descriptors, allowing downstream statistical modeling and domain-specific interpretation to be conducted as appropriate for individual study designs. This design choice reflects the exploratory and hypothesis-generating role of spatial feature analysis in computational pathology, particularly in settings where spatial heterogeneity and sampling variability are inherent. Complete list of visual examples by significant feature can be found in the supplementary document.
5.4 Sensitivity of Exploratory Findings to Significance Thresholds
In the NLST dataset, $16$ spatial features reached significance at $p<0.20$, with $11$ features remaining significant at $p<0.15$, 77 features at $p<0.10$, and $3$ features at $p<0.05$. Similarly, the OPMD dataset demonstrated $15$ significant features at $p<0.20$, $10$ features at $p<0.15$, 88 features at $p<0.10$, and $3$ features achieving significance at $p<0.05$. This monotonic reduction in the number of associated features across increasingly stringent thresholds reflects the sensitivity of exploratory spatial findings to threshold choice, rather than providing evidence of optimal cutoff values. These observations indicate that spatial descriptors may exhibit moderate and context-dependent associations with clinical outcomes, which is consistent with the inherent variability of cellular spatial arrangements in tumor tissue architecture and heterogeneity across disease sites.
@@ 5.4 Sensitivity of Exploratory Findings to Significance Thresholds
The use of multiple significance thresholds is justified by several factors inherent to spatial point pattern analysis. First, the resolution and density of marked point pattern data can influence the sensitivity of spatial statistics to detect biologically relevant structures. Second, the complex multi-scale nature of tumor spatial organization may yield moderate rather than strong statistical associations. Third, given the exploratory role of comprehensive spatial feature analysis, considering a range of nominal significance thresholds allows potentially informative spatial patterns to be identified for further investigation in larger cohorts or independent validation studies, rather than being prematurely excluded by a single stringent cutoff.
@@ 8 Data Availability Statement
Case study II. The OPMD dataset analyzed in this study is available in the Zenodo repository under restricted access at https://doi.org/10.5281/ZENODO.10664056. This dataset is described in (Piyarathne et al., 2024). The nuclei segmentation and classification results generated in this study using ConvPath (Wang et al., 2019) are available from the corresponding author upon reasonable request.